OpenStudy (anonymous):

What is the solution to G over 2 = 1 over 5 ? 0.1 0.4 4 10

hero (hero):

Hint: Cross Multiply to get G(5) = 1(2)

OpenStudy (anonymous):

And then convert the fraction that you get into a decimal.

OpenStudy (anonymous):

So we have G/2=1/5. What we can do here is to cross-multiply 2 to both sides like hero said.  Now we have G=2/5 which is 0.4, answer B. to convert a fraction to a decimal just divide 2 into 5 and you'll get 0.4 if you didn't know that :) -Hope it helps :)